If 6.77 moles of an ideal gas has a pressure of 1.61 atm, and a volume of 29.75 L, what is the temperature of the sample?
Answer #1

As we know that the ideal gas equation as

pV=nRT

Here p= pressure = 1.61 atm

V= 29.75 L

R = gas constant =0.08LatmK^{-1}mol^{-1}

T= temperature in Kelvin

n = moles = 6.77 mol

Now apply

pV=nRT

\\1.61atm\times 29.75L=6.77 mol\times 0.08LatmK^{-1}mol^{-1}\times T\\ 47.9Latm=0.5416\times T\\ T=\frac{47.9}{0.5416}K\\ T=88.44K

Hence the temperature is 88.44 K